OpenStudy (anonymous):

A projectile is shot from the edge of a cliff 125 m above the ground. It has an initial velocity 50.0 m/s 37.0° above the horizontal at the launch point. a) Determine the time taken to reach the ground. b) How far is this point from the base of the cliff? c) What is the magnitude and direction of the final velocity just before the point of impact?

OpenStudy (anonymous):

was there an equation given

OpenStudy (anonymous):

no but its based off the three equations of constant acceleration so, s=vit + 1/2at^2 vf=vi + at vf^2=vi^2 + 2as

OpenStudy (anonymous):

you have no idea how to do it? atleast to start with?

OpenStudy (anonymous):